# Completeness conditions for spacetimes with low-regularity metrics

### v1

@@ Abstract.
We extend Beem’s three completeness notions—finite compactness, timelike Cauchy completeness, and Condition A —originally defined for spacetimes, to Lorentzian length spaces and study their relationships. We prove that finite compactness implies timelike Cauchy completeness and that timelike Cauchy completeness implies Condition A for globally hyperbolic Lorentzian length spaces. Furthermore, for globally hyperbolic $C^{1}$-spacetimes under suitable assumptions ensuring good behavior of causal geodesics, we show that the three conditions are equivalent. These results can be regarded as a Hopf-Rinow type theorem for low-regularity Lorentzian geometry. The appendix presents examples of $C^{1}$-spacetimes—where geodesic uniqueness may fail— in which causal geodesics nevertheless behave well, illustrating the scope of our results.
@@ 1. Introduction
Beem [B76] introduced finite compactness and timelike Cauchy completeness, which are Lorentzian analogues of bounded compactness and metric completeness for Riemannian manifolds, respectively. Beem proved a following Lorentzian version of the Hopf–Rinow theorem for $C^{2}$-spacetimes. Here, a $C^{2}$-spacetime means a smooth manifold endowed with a Lorentzian metric of class $C^{2}$.
@@
Therefore, the question of completeness remains poorly understood for spacetimes with Lorentzian metrics of low-regularity.
@@
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and proper $C^{1}$-spacetime such that the exponential map as defined in Definition 4.8 is continuous. Then the following three conditions are equivalent:
@@ Acknowledgements.
The author would like to thank Kotaro Yamada for helpful discussions and valuable advice. This work was supported by JST SPRING, Japan Grant Number JPMJSP2180.
@@ 2. Preliminaries
To avoid some technical difficulties, we impose the following three conditions, which are used in the definition of Lorentzian length spaces.
@@
Definition 2.9 ([KS18, Definition 3.4]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and $x\in X.$ A neighborhood $U$ of $x$ is called causally closed if for any sequences $(p_{n})_{n},(q_{n})_{n}$ in $U$ satisfying $p_{n}\leq q_{n}$ for all $n\in\mathbf{N}$ and $p_{n}\to p\in\overline{U},\ q_{n}\to q\in\overline{U}$, then $p\leq q$, where $\overline{U}$ is the closure of $U.$ A Lorentzian pre-length space $(X,\ll,\leq,d,\tau)$ is called locally causally closed if each point in $X$ has a causally closed neighborhood.
@@
A Lorentzian length space is a causally path connected, locally causally closed and localizable Lorentzian pre-length space satisfying
@@
A fundamental source of Lorentzian length spaces comes from spacetimes with low-regularity Lorentzian metrics. However, one cannot go below Lipschitz regularity in general: for metrics of regularity strictly less than $C^{0,1}$, there exist spacetimes that fail even to be Lorentzian pre-length spaces (cf. [KS18, Example 5.2]).
@@
To avoid pathological examples, we need to impose some causal conditions. We now introduce the following conditions.
@@ 3. Completeness conditions and their implications
The following definition is a Lorentzian analogue of bounded compactness. Beem [B76] first introduced it for spacetimes.
@@
The following definition is a Lorentzian analogue of metric completeness. Beem [B76] first introduced it for spacetimes.
@@ 4. Geodesics in $C^{1}$-spacetimes
In a $C^{1}$-spacetime, the geodesic equation is merely continuous, so the geodesic ODE has locally solutions but uniqueness may fail. Indeed, [KOSS22, Example 3.2] exhibits a $C^{1}$-spacetime in which maximal timelike geodesics and null geodesics branch. Consequently, to guarantee the uniqueness of causal geodesics, we impose a non-branching condition on causal geodesics (cf. [KOSS22, Definition 3.1]). We also show that this non-branching hypothesis is actually satisfied by a class of $C^{1}$ low-regularity spacetimes treated in Appendix A; see Theorem A.3.
@@
We define the following sufficient condition ensuring that maximal causal geodesics do not branch.
@@
Let $(M,g)$ be a $C^{1}$-spacetime and $p\in M$. We say that $(M,g)$ is proper at $p$ if there exists a neighborhood $U$ of $p$ such that, for any two distinct maximal causal geodesics $\gamma_{1},\gamma_{2}$ starting at $p$, the intersection set $\gamma_{1}\cap\gamma_{2}\cap U$ has no accumulation point in $U$. We say $(M,g)$ is proper if it is proper at every $p\in M$.
@@
Let $(M,g)$ be a causally non-branching $C^{1}$-spacetime, proper at $p\in M$ in the sense of Definition 4.6. Suppose $\gamma_{1},\gamma_{2}:[0,1]\to M$ are maximal causal geodesics with
@@
Assume $\gamma_{1}\neq\gamma_{2}$. By properness at $p$, there is $\varepsilon>0$ with $\gamma_{1}((0,\varepsilon])\cap\gamma_{2}((0,\varepsilon])=\emptyset$. Let $\gamma_{-}\colon(-\delta,0]\to M$ be a past-directed causal geodesic with $\gamma_{-}(0)=p$ and $\dot{\gamma}_{-}(0)=v$, obtained from local existence for the geodesic ODE. Define concatenations
(4.5) $\tilde{\gamma}_{i}(t)=\begin{cases}\gamma_{-}(t),&t\in(-\delta,0],\\ \gamma_{i}(t),&t\in[0,\varepsilon],\end{cases}\qquad i=1,2.$
Each $\tilde{\gamma}_{i}$ is a causal geodesic on $(-\delta,\varepsilon]$, the two coincide on $(-\delta,0]$, but they differ on $(0,\varepsilon]$. Thus we have a causal geodesic that branches at the interior point $t=0$, contradicting causal non-branching. Hence $\gamma_{1}=\gamma_{2}$ on $0,\varepsilon]$. By [KOSS22, Lemma 3.3], they coincide on the whole $[0,1]$. ∎
Since the maximal geodesic does not branch in the above situation, we can define the exponential map as follows.
@@
Let $(M,g)$ be a causally non-branching, proper $C^{1}$-spacetime and let $p\in M$. Then, we can define the domain of the exponential map at $p$ by
(4.6) $\mathcal{E}_{p}\coloneqq\{v\in T_{p}M;\text{$\gamma_{v}$ is defined on an interval containing $[0,1]$}\}$
and the exponential map $\exp_{p}\colon\mathcal{E}_{p}\to M$ at $p$ by
@@
where $\gamma_{v}$ is the maximal causal geodesic with $\gamma(0)=p$ and $\dot{\gamma}(0)=v$.
@@ A $C^{1}$-spacetime is called globally hyperbolic if
Finally, we show that condition A implies finite compactness for $C^{1}$-spacetime if the exponential map is continuous.
@@
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and proper $C^{1}$-spacetime such that the exponential map as defined in Definition 4.8 is continuous. If $(M,g)$ satisfies Condition A, then $(M,g)$ is finitely compact.
@@
Let $\lVert\ \rVert$ be the Euclidean inner product on $T_{q}M$. For each direction $L$ at $q$ let $u(L)$ be the unique future-pointing causal vector in $L$ with $\lVert u(L)\rVert=1$. Put $u_{n}\coloneqq u(L_{n})$ and $u_{0}\coloneqq u(L_{0})$. Since $L_{n}\to L_{0}\ (n\to\infty)$, we have $u_{n}\to u_{0}\ (n\to\infty)$. For each $n$ write
@@
By continuity of $\exp_{q}$ and of $T$, the map $F$ is continuous in $(u,s)$ on the domain of exponential map and by Lemma 4.10, for each fixed $u$ the function $s\mapsto F(u,s)$ is strictly increasing.
@@
The continuity of $F$ and $u_{n}\to u_{0}$ yield $F(u_{n},s_{0})\to F(u_{0},s_{0})=b\ (n\to\infty)$, whereas we also have $F(u_{n},s_{n})\to b$ as $n\to\infty$—a contradiction. Hence $s_{n}\to s_{0}$, and with $u_{n}\to u_{0}$ we conclude $X_{n}=s_{n}u_{n}\to s_{0}u_{0}=X_{0}$
Finally, by continuity of $\exp_{q}$ we obtain $z_{n}=\exp_{q}(X_{n})\to\exp_{q}(X_{0})=x(t_{0})\in K_{1}$. Therefore every sequence in $K_{1}$ has a convergent subsequence with limit in $K_{1}$, so $K_{1}$ is compact. This proves finite compactness. ∎
@@
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and proper $C^{1}$-spacetime such that the exponential map as defined in Definition 4.8 is continuous. Then the following three conditions are equivalent:

### v2

@@ Abstract.
We extend Beem’s three completeness notions — finite compactness, timelike Cauchy completeness, and Condition A — originally defined for spacetimes, to Lorentzian length spaces and study their relationships. We prove that finite compactness implies timelike Cauchy completeness and that timelike Cauchy completeness implies Condition A for globally hyperbolic Lorentzian length spaces. Furthermore, for globally hyperbolic $C^{1}$-spacetimes, we establish the equivalence of the three conditions assuming the causally non-branching and non-intertwining conditions, which in fact imply the continuity of the causal exponential map. These results can be regarded as a Hopf-Rinow type theorem for low-regularity Lorentzian geometry. The appendix presents examples of $C^{1}$-spacetimes — where geodesic uniqueness may fail — in which causal geodesics nevertheless behave well, illustrating the scope of our results.
@@ 1. Introduction
Busemann [B67] first introduced finite compactness and timelike Cauchy completeness for timelike spaces and proved that finite compactness implies timelike Cauchy completeness. These conditions are Lorentzian analogues of bounded compactness and metric completeness for Riemannian manifolds, respectively.
Although the specific terminology of timelike spaces introduced by Busemann may not appear frequently in standard Lorentzian texts, his approach has found interest in modern frameworks. Indeed, the structures introduced by Busemann are fundamental to currently active subjects of research, particularly concerning the study of convex subsets and their duality in Lorentzian geometries (cf. [FS19]).
Building upon Busemann’s work, Beem [B76] proved a following Lorentzian version of the Hopf–Rinow theorem for $C^{2}$-spacetimes. Here, a $C^{2}$-spacetime means a smooth manifold endowed with a time-orientable Lorentzian metric of class $C^{2}$.
@@
Therefore, the question of completeness remains insufficiently understood for spacetimes with Lorentzian metrics of low-regularity. Extending Hopf–Rinow type theorems to such low-regularity settings is of intrinsic mathematical interest. For instance, Buro [B23, Theorem 5.22] established a Hopf–Rinow type theorem for weak Finsler metrics of low regularity by introducing the notion of special geodesics.
@@
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and non-intertwining $C^{1}$-spacetime. Then the following three conditions are equivalent:
@@ Acknowledgements.
The author would like to thank Sumio Yamada and Kotaro Yamada for helpful discussions and valuable advice. The author would also like to thank the two anonymous reviewers for their careful reading and constructive comments. In particular, the author is grateful for their suggestion that helped to weaken the assumptions of the main theorem. This work was supported by JST SPRING, Grant Number JPMJSP2180.
@@ 2. Preliminaries
To avoid some technical difficulties, we impose the following three conditions, which are used in the definition of Lorentzian length spaces. We note that the notion of local causal closedness defined in [KS18], which is included in the definition of Lorentzian length spaces, is not appropriate for smooth spacetimes that are not strongly causal (cf. [ACS20, Example 2.20]), whereas weak local causal closedness defined in [ACS20, Definition 2.19] is suited. However, since we assume global hyperbolicity for Lorentzian length spaces throughout this article, the two notions are equivalent (cf. [ACS20, Proposition 2.21]); thus, it does not matter which definition we adopt.
@@
Definition 2.9 ([ACS20, Definition 2.19]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and $x\in X.$ A neighborhood $U$ of $x$ is called weakly causally closed if for any sequences $(p_{n})_{n},(q_{n})_{n}$ in $U$ satisfying $p_{n}\leq_{U}q_{n}$ for all $n\in\mathbf{N}$ and $p_{n}\to p\in U,\ q_{n}\to q\in U$, then $p\leq_{U}q$. A Lorentzian pre-length space $(X,\ll,\leq,d,\tau)$ is called locally weakly causally closed if each point in $X$ has a weakly causally closed neighborhood.
@@
A Lorentzian length space is a causally path connected, locally causally closed, and localizable Lorentzian pre-length space satisfying
@@
A fundamental source of Lorentzian length spaces comes from spacetimes with low-regularity Lorentzian metrics. We note that any $C^{0}$-spacetime can be formulated as a Lorentzian pre-length space by using nearly timelike curves [L25, Corollary 3.6]. However, spacetimes with regularity strictly less than $C^{0,1}$ may fail to be Lorentzian length spaces due to the lack of (weakly) local causal closedness [L25, Example 3.7].
@@
To avoid pathological examples, we need to impose some causality conditions. We now introduce the following conditions.
@@ 3. Completeness conditions and their implications
The following definition is a Lorentzian analogue of bounded compactness. Busemann [B67] first introduced it for timelike spaces.
@@
The following definition is a Lorentzian analogue of metric completeness. Busemann [B67] first introduced it for timelike spaces.
@@
Busemann [B67] first proved the following result for $C^{2}$-spacetimes. We extend this result to globally hyperbolic Lorentzian length spaces.
@@ 4. Geodesics in $C^{1}$-spacetimes
In a $C^{1}$-spacetime, the geodesic equation is merely continuous, so the geodesic ODE has local solutions, but uniqueness may fail. Indeed, [KOSS22, Example 3.2] exhibits a $C^{1}$-spacetime in which maximal timelike geodesics and null geodesics branch. Consequently, to guarantee the uniqueness of causal geodesics, we impose causally non-branching which is a stronger condition than maximally causally non-branching (cf. [KOSS22, Definition 3.1]). We also show that this non-branching hypothesis is actually satisfied by a class of $C^{1}$ low-regularity spacetimes treated in Appendix A; see Theorem A.2.
@@
We define the following sufficient condition ensuring that causal geodesics do not branch.
@@
Let $(M,g)$ be a $C^{1}$-spacetime and $p\in M$. We say that $(M,g)$ is non-intertwining at $p$ if there exists a neighborhood $U$ of $p$ such that, for any $\varepsilon>0$ and any causal geodesics $\gamma_{1}\colon[0,\varepsilon]\to U$ and $\gamma_{2}\colon[0,\varepsilon]\to U$ starting at $p$ with the same initial velocity $v$, it holds that $\gamma_{1}(t)=\gamma_{2}(t)$ for all $t\in[0,\varepsilon]$ or $\gamma_{1}((0,\varepsilon])\cap\gamma_{2}((0,\varepsilon])$ has no accumulation point. We say $(M,g)$ is non-intertwining if it is non-intertwining at every $p\in M$.
The causally non-branching and non-intertwining conditions ensure that any causal geodesic with the same initial condition is unique.
@@
Let $(M,g)$ be a causally non-branching and non-intertwining $C^{1}$-spacetime. Suppose $\gamma_{1},\gamma_{2}\colon[0,1]\to M$ are causal geodesics with the same initial condition:
@@
First, we show that $\gamma_{1}$ and $\gamma_{2}$ coincide on a small interval $0,t_{1}]$ where $t_{1}>0$. Suppose, for the sake of contradiction, that $\gamma_{1}$ and $\gamma_{2}$ do not coincide on any neighborhood of 0. The non-intertwining condition implies that there exists $\varepsilon>0$ such that $\gamma_{1}((0,\varepsilon])\cap\gamma_{2}((0,\varepsilon])$ has no accumulation point. This guarantees the existence of a possibly smaller $\eta\in(0,\varepsilon]$ such that $\gamma_{1}(t)\neq\gamma_{2}(t)$ for all $t\in(0,\eta]$.
Let $\gamma_{-}\colon(-\delta,0]\to M$ be a past-directed causal geodesic with $\gamma_{-}(0)=p$ and $\dot{\gamma}_{-}(0)=v$, obtained from local existence for the geodesic ODE. Define concatenations
(4.5) $\tilde{\gamma}_{i}(t)=\begin{cases}\gamma_{-}(t),&t\in(-\delta,0],\\ \gamma_{i}(t),&t\in[0,\eta],\end{cases}\qquad i=1,2.$
Each $\tilde{\gamma}_{i}$ is a causal geodesic on $(-\delta,\eta]$, the two coincide on $(-\delta,0]$, but they differ on $(0,\eta]$. Thus, we have a causal geodesic that branches at the interior point $t=0$, contradicting causal non-branching.
Therefore, the assumption is false, and $\gamma_{1}$ and $\gamma_{2}$ must coincide on some small interval. By a standard connectedness argument, this local uniqueness extends to the whole interval $[0,1]$. ∎
Since the causal geodesic with a given initial condition is unique in the above situation, we can define the causal exponential map as follows.
@@
Let $(M,g)$ be a causally non-branching and non-intertwining $C^{1}$-spacetime and let $p\in M$. Then, we can define the domain of the causal exponential map at $p$ by
(4.6) $\mathcal{E}_{p}\coloneqq\{v\in T_{p}M;g_{p}(v,v)\leq 0,\text{$\gamma_{v}$ is defined on an interval containing $[0,1]$}\}$
and the causal exponential map $\exp_{p}\colon\mathcal{E}_{p}\to M$ at $p$ by
@@
where $\gamma_{v}$ is the causal geodesic with $\gamma(0)=p$ and $\dot{\gamma}(0)=v$.
It is worth noting that for $C^{1,1}$-spacetimes, the exponential map is a bi-Lipschitz homeomorphism locally around any point [KSS14, Theorem 2.1]. On the other hand, for the $C^{1}$-spacetimes considered here, we establish the continuity of the causal exponential map. This follows from standard ODE theory, and the proof proceeds along the lines of [KOSS22, Corollary 2.6], adapted to the present context.
Proposition 4.9.
Let $(M,g)$ be a causally non-branching, non-intertwining $C^{1}$-spacetime. Then the causal exponential map $\exp_{p}\colon\mathcal{E}_{p}\to M$ at $p$ is continuous for every $p\in M$.
Proof.
It suffices to consider the case where the geodesic is contained in a single coordinate chart. The result extends to the general case by a finite covering argument, noting that the convergence of position and velocity at the end of one chart implies the convergence of initial data for the next.
Let $\{v_{k}\}_{k\in\mathbf{N}}\subset\mathcal{E}_{p}$ be a sequence of causal vectors converging to $v\in\mathcal{E}_{p}$. In a local coordinate $(U,\phi)$, the geodesic equation with initial point $\tilde{p}=(\tilde{p}^{1},\dots,\tilde{p}^{n})$ and initial velocity $\tilde{v}=(\tilde{v}^{1},\dots,\tilde{v}^{n})$ is rewritten as first order differential equations. Let $\tilde{y}(t)=(\tilde{y}^{1}(t),\dots,\tilde{y}^{2n}(t))^{T}$. The equation is as follows:
(*) $\frac{d}{dt}\begin{pmatrix}\tilde{y}^{1}(t)\\ \vdots\\ \tilde{y}^{n}(t)\\ \tilde{y}^{n+1}(t)\\ \vdots\\ \tilde{y}^{2n}(t)\end{pmatrix}=\begin{pmatrix}\tilde{y}^{n+1}(t)\\ \vdots\\ \tilde{y}^{2n}(t)\\ -\Gamma^{1}_{ij}(\tilde{y}^{1},\dots,\tilde{y}^{n})\tilde{y}^{n+i}\tilde{y}^{n+j}\\ \vdots\\ -\Gamma^{n}_{ij}(\tilde{y}^{1},\dots,\tilde{y}^{n})\tilde{y}^{n+i}\tilde{y}^{n+j}\end{pmatrix}$
with initial conditions:
$\tilde{y}^{l}(0)$ $=\tilde{p}^{l}\quad(1\leq l\leq n),$
$\tilde{y}^{l}(0)$ $=\tilde{v}^{l-n}\quad(n+1\leq l\leq 2n).$
Define the function $F$ by the right-hand side of (*4):
[…]
Then $F$ is continuous since the metric $g$ is $C^{1}$.
Let $Y_{k}$ be the solution of (*4) with the initial conditions corresponding to $p=(p^{1},\dots,p^{n})$ and $v_{k}=(v_{k}^{1},\dots,v_{k}^{n})$. Since $v_{k}\to v$, the initial data converges as $(p,v_{k})\to(p,v)$. By [KOSS22, Proposition 2.5], for any subsequence of $\{Y_{k}\}$, there exists a further subsequence $\{Y_{k_{m}}\}$ and a solution $\tilde{Y}$ of (*4) with the initial condition corresponding to $(p,v)$ such that
$Y_{k_{m}}\to\tilde{Y}\qquad\text{in }C^{1}([0,1],\mathbf{R}^{2n}).$
Let $Y$ be the solution of (*4) corresponding to $p=(p^{1},\dots,p^{n})$ and $v=(v^{1},\dots,v^{n})$. By Lemma 4.7, any two causal geodesics with the same initial conditions coincide. Therefore, the limit solution $\tilde{Y}$ obtained from the subsequence must coincide with the solution $Y$, which defines the causal exponential map $\exp_{p}$.
Since every convergent subsequence converges to the same unique limit $Y$, the entire sequence $\{Y_{k}\}$ converges to $Y$ in $C^{1}([0,1],\mathbf{R}^{2n})$. In particular, evaluating at $t=1$, the position components converge:
$\lim_{k\to\infty}(Y_{k}^{1}(1),\dots,Y_{k}^{n}(1))=(Y^{1}(1),\dots,Y^{n}(1)).$
This means $\lim_{k\to\infty}\exp_{p}(v_{k})=\exp_{p}(v)$, proving the continuity of the exponential map. ∎
@@ A $C^{1}$-spacetime is called globally hyperbolic if
Finally, we show that Condition A implies finite compactness for a $C^{1}$-spacetime.
@@
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and non-intertwining $C^{1}$-spacetime. If $(M,g)$ satisfies Condition A, then $(M,g)$ is finitely compact.
@@
Let $\lVert\ \rVert$ be a Euclidean inner product on $T_{q}M$. For each direction $L$ at $q$ let $u(L)$ be the unique future-pointing causal vector in $L$ with $\lVert u(L)\rVert=1$. Put $u_{n}\coloneqq u(L_{n})$ and $u_{0}\coloneqq u(L_{0})$. Since $L_{n}\to L_{0}\ (n\to\infty)$, we have $u_{n}\to u_{0}\ (n\to\infty)$. For each $n$ write
@@
By Proposition 4.9 and the continuity of $T$, the map $F$ is continuous in $(u,s)$ on the domain of exponential map and by Lemma 4.11, for each fixed $u$ the function $s\mapsto F(u,s)$ is strictly increasing.
@@
The continuity of $F$ and $u_{n}\to u_{0}$ yield $F(u_{n},s_{0})\to F(u_{0},s_{0})=b\ (n\to\infty)$, whereas we also have $F(u_{n},s_{n})\to b$ as $n\to\infty$, a contradiction. Hence $s_{n}\to s_{0}$, and with $u_{n}\to u_{0}$ we conclude $X_{n}=s_{n}u_{n}\to s_{0}u_{0}=X_{0}$.
Finally, by Proposition 4.9, we obtain $z_{n}=\exp_{q}(X_{n})\to\exp_{q}(X_{0})=x(t_{0})\in K_{1}$. Therefore every sequence in $K_{1}$ has a convergent subsequence with limit in $K_{1}$, so $K_{1}$ is compact. This proves finite compactness. ∎
@@
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and non-intertwining $C^{1}$-spacetime. Then the following three conditions are equivalent: